Alma Cataloging User Group - August 24, 2020
August 24, 2020 – Academic Cataloging Group Notes
ATTENDING: Amy (NDSCS), Aubrey (NMY), Ben (NDV), Drew (UND), Erika (UNF), Ginny (ODIN), Felecia (UND), Jasmine (NDJ), Jason (ODIN), Laurie (UNE), Liz (NBJ), Megan (ODIN), Phyllis Bratton, Staci (NDI), Tina (NDSU), Lynn (ODIN)
Data Cleanup/Data Review:
- UND - browsing call numbers, UND - cutters used in 099 tags, NMI -- cutters with 092 for children’s collections, virtual browse uses 050 instead of 090 (based on bib record rather than holding) – problem listed in Basecamp.
- Children’s fiction call numbers are often not in order (TD12709890/SF00821559). UPDATE: Case sent to development issue should be fixed in the September 2020 release. Deletes/updates processing – UPDATE: Configuration fixed. Further testing needed.
- UND has trouble matching with brief order records in the IZ – Felecia sent examples – waiting on workflows.
- Shelby sent email 8/5 - norm rule delete undesirable fields from OCLC records that are imported daily for discussion. Last week discussions included 653 being used. 690 field for notes and ND author but 691 should be ND Author. Further discussion? Please review.
- List of reports to demo in Analytics – PLEASE add to Basecamp post https://3.basecamp.com/3765443/buckets/14394235/documents/2942042253 - Premium sandbox – need user credentials assigned and then ready to roll out.
- NetLibrary/EBSCOeBooks problem for several libraries. Need further investigation to define issue.
- OCLC ODINA daily extracts –Revisit workflow.
Note: Info on https://www.orbiscascade.org/alma-daily-bib-record-updates/:
Due to OCLC file delivery and Alma import profile schedules, records edited in Connexion will be updated in Alma two business days after the Connexion update. If immediate updating of a bib record in Alma is desired, release the record from the Alma Metadata Editor and export (F5) the record from Connexion to the Network Zone.
- Bib importing not bringing in updated record from OCLC (NDSU reported). More information is being gathered and tracked. https://3.basecamp.com/3765443/buckets/14394235/messages/2950835911#__r…
- Overdrive access issue – NMY – SIP servers. ODIN Office working on temporary fix via ALEPH.